a convex lens and a concave lens with equal focal length are cemented together and exposed to a beam of light falling on the combination symmetrically with respect to optical axes. find the distance aT WHICH THE BEAM WILL COME TO A frocus on the other side of the lens.

Grade:10

1 Answers

vikas askiitian expert
13 years ago

combined focal length = f

1/f = 1/f1 + 1/f2 - d/f1f2

when lences are kept very close then d =0

1/f = 1/f1 + 1/f2

f1 = -F  (concave)

f2 = F (convex)

1/f = 0

f = infinity

so when parallel beam of light is incident then image is formed at focus = infinity ..